OpenStudy (anonymous):

what is the area of rectangle L=8, W=4

OpenStudy (gorv):

area = L*W

OpenStudy (ahsome):

\(Area=l*w\) Subsitute \(l\) and \(w\) \(Area=l*w\) \(Area=8*4\) \(Area=32\)

OpenStudy (itsbribro):

l*w= length x width 8*4= area or a 8*4= 32